OpenStudy (blake57roger):

For the past 5 months Owen's lawn care business had these profits or losses: $110, –$75, $80, –$50, and –$25. What was his average profit or loss for these five months

OpenStudy (blake57roger):

A. His average loss was $68 per month. B. His average loss was $8 per month. C. His average profit was $8 per month. D. His average profit was $68 per month.

OpenStudy (blake57roger):

@gorv

OpenStudy (gorv):

negative sign means here is loss total loss = 75+ 50 + 25=150 profit =110+80=190 so overall profit=190-150=40but it is for 5 month so for one month average=40/5=8

OpenStudy (blake57roger):

thanks